OpenStudy (anonymous):

is the derivative e^x^2 the same as the derivative e^2x? why doesnt this hold?

hartnn (hartnn):

simply because x^2 is NOT = 2x those 2 are different function, hence their derivative will also be different! :)

OpenStudy (ikram002p):

if you mean e^(x^2) and e^(2x) then no but e^x^2=e^(2x)

hartnn (hartnn):

\((e^x)^2 = e^{2x} \)

OpenStudy (loser66):

Let \(t= e^x\) so that if it is \((e^x)^2\), then it is \(t^2\) if it is \(e^{2x} = e^{x+x} = e^x*e^x= t*t =t^2\) So that I don't see why they are not the same @ikram002p

hartnn (hartnn):

(e^x)^2 is indeed e^(2x) but e^x^2 is not e^(2x) infact, e^x^2 is ambiguous :P

geerky42 (geerky42):

a^b^c^b^e = \(a^{b^{c^{d^{e}}}} = a^{\left(b^{\left(c^{(d^{e})}\right)}\right)}\neq \left(\left((a^b)^c\right)^d\right)^e = a^{bcde}\) So e^x^2 \(= e^{x^2}\neq(e^x)^2 = e^{2x}\)

OpenStudy (anonymous):

this was the whole question: determine the derivative of the function y(x) = ((e^x)^2) ln (2x + 0.5) in x=0 for the derivative of ((e^x)^2), the right answer was 2x. ((e^x)^2) but if i turned ((e^x)^2) into e^2x then the derivative would have been just 2.e^2x.

hartnn (hartnn):

"the derivative of ((e^x)^2), the right answer was 2x. ((e^x)^2) " Nopes. (e^x)^2 does equal e^(2x) and its derivative is 2 e^(2x) and not 2x (e^x)^2

hartnn (hartnn):

the derivative of e^(x^2) will be 2x e^(x^2)

hartnn (hartnn):

and e^(x^2) is not in your question

OpenStudy (anonymous):

there are no brackets in the book so i'm kind a confused, it's 'e' powered to the x powered to 2

OpenStudy (anonymous):

slope of two different things would different use simple logic